OpenStudy (anonymous):

Josephine develops her own rules for playing the game of poker. One of her rules is that you can have a hand called a “jojo” where all 5 cards are the same color. The probability you are dealt a “jojo” from a standard 52 card playing deck can be written in the form \[\frac{j \times _{26}C_{l}}{_{52}C_{n}}\] What is the value of j + l + n ?

OpenStudy (anonymous):

\[\frac{j \times _{26}C_{l}}{_{52}C_{n}} = \frac{j \times \frac{26!}{(26 - l)!l!}}{\frac{52!}{(52 - n)!n!}} = j \times \frac{26!}{(26 - l)!l!} \times \frac{(52 - n)!n!}{52!}\] I got this far. What should I do now? @Wired

OpenStudy (phi):

I would not expand it. I would try to figure out the probability of getting 5 cards of the same color from a standard deck the denominator should be the number of hands with 5 cards. the top is the number of hands with 5 cards of the same color.

OpenStudy (phi):

order does not matter for a "jojo" so I would say 52C5 are the number of hands so you know n=5 the number of red hands would be 26C5 same for black hands so l=5 j=2 sum= 12

OpenStudy (anonymous):

\[_{52}C_{5}\] = (52*51*50*49*48)/5! = 10*2*49*51*52 Is that for the denominator?

OpenStudy (phi):

yes, but the actual value is not important here.

OpenStudy (phi):

unless you are learning what M CHOOSE N means

OpenStudy (phi):

see my post up above

OpenStudy (anonymous):

THanks! I didn't think of doing that :)